Medicare would like to test the hypothesis that the average monthly rate for one-bedroom assisted-living facility is equal to $3,300. A random sample of 12 assisted-living facilities had an average rate of $3,690 per month. The standard deviation for this sample was $530. Medicare would like to set α =  0.05. The test statistic for this hypothesis test would be ________.

1.98

2.55

-2.16

-1.37
Answer #1

Solution :

This is the two tailed test,  

The null and alternative hypothesis is ,

H0 :  \mu = 3300

Ha : \mu   \neq 3300

Test statistic = z

= (T - \mu ) / \sigma / \sqrt n

= (3690 - 3300) / 530 / \sqrt 12

Test statistic = z = 2.55